an electron (me = 9.10939 × 10–28 g) moving at 4.00 × 106 m/s in an electron microscope


1
Expert's answer
2021-04-07T11:37:15-0400

"λ = \frac{h}{p} = \frac{h }{ m × v}"


"\lambda=\frac{6.62 × 10 ^{-34} }{ [(9.10939 ×10^{-28 })×(4.0×10^6)]}"


"\lambda= 1.816806614×10^{-13} m"


"\lambda=0.0001816806614 nm"